OpenStudy (anonymous):

A traffic light follows the pattern green, yellow, red for 60, 5, and 20 seconds, respectively. What is the probability that a driver approaching this light will find it green or yellow?

OpenStudy (anonymous):

\[\frac{60+5}{60+5+20} = \frac{65}{85} = \frac{13}{17}\]